Probability Theory and Geometric Analysis

確率論と幾何解析

科学研究費補助金基盤 (A) (研究代表者：重川一郎谷口説男) による上記シンポジウムを開催します。

日時： 2003 年 9 月 16 日 (火)13:30〜9 月 18 日 (木)17:20 場所：横浜市立大学金沢八景キャンパスカメリアホール

(横浜市金沢区瀬戸 22-2 京浜急行金沢八景駅下車徒歩１０分)

世話人：桑江一洋横浜市立大学総合理学研究科 (kuwae@yokohama-cu.ac.jp)
       小倉幸雄佐賀大学理工学部 (ogura@ms.saga-u.ac.jp)

Program

16 September

13:30–14:20 Mark Pinsky (Northwestern University)

Local stochastic differential geometry:

feeling the shape of a manifold with Brownian motion

14:30–15:20 Satoshi Ishiwata (Tohoku University) 石渡聡 (東北大学) Some analytic properties on nilpotent covering graphs

15:30–16:20 Motoko Kotani (Tohoku University) 小谷元子 (東北大学) The magnetic transition operators on a crystal lattice

16:30–17:20 Atsushi Katsuda (Okayama University) 勝田篤 (岡山大学) Heat kernels on nilpotent coverings and related topics

17 September

10:00–10:50 Junya Takahashi (Tohoku University) 高橋淳也 (東北大学)

Vanishing of cohomology groups and large eigenvalues of the Laplacian on p-forms

11:10–12:00 Atsushi Kasue (Kanazawa University) 加須栄篤 (金沢大学) Convergence of Riemannian manifolds and harmonic maps

13:30–14:20 Shin-ichi Ohta (Tohoku University) 太田慎一 (東北大学) Regularity of harmonic functions on metric spaces 14:30–15:20 Takefumi Kondo (Kyoto University) 近藤剛史 (京都大学)

Probability distribution of metric measure spaces

15:30–16:20 Takashi Shioya (Tohoku University) 塩谷隆 (東北大学)

Geometric aspects of variational convergence I

16:30–17:20 Short communications

18 September

10:00–10:50 Yusuke Higuchi (Showa University) 樋口雄介 (昭和大学)

Non-bipartiteness of graphs and the upper bounds of Dirichlet forms

11:00–11:50 Takashi Shioya (Tohoku University) 塩谷隆 (東北大学)

Geometric aspects of variational convergence II

13:30–14:20 Mark Pinsky (Northwestern University)

Pointwise Fourier inversion and related eigenfunction expansions

14:30–15:20 Kazumasa Kuwada (Kyoto University) 桑田和正 (京都大学)

On large deviations for random currents induced from stochastic line integrals

15:30–16:20 Hiroshige Shiga (Tokyo Institute of Technology) 志賀啓成 (東京工業大学) Deformation of Riemann surfaces and potential theory

16:30–17:20 Toshikazu Sunada (Meiji University) 砂田利一 (明治大学)

Crystal lattices and convex polyhedra

注意 1 ：横浜市立大学金沢八景キャンパスへのアクセスは http://www.yokohama-cu.ac.jp/access/access.html を参照してみてください。京浜急行金沢八景駅へは、横浜駅から快速特急 (快特) で金沢文庫駅まで１４分程で、そこから乗り換えして普通列車で２分程です。金沢文庫駅で乗り換えを忘れると快特は横須賀中央駅まで止まりませんのでご注意ください。

注意２：大学の近辺はホテルが僅少です。ホテルニッコー金沢八景 (JAL とは無関係) がありますが客層は釣り舟乗船者が主体です。ホテルニッコー金沢八景〒 236-0028 横浜市金沢区州崎町 5-50 Tel 045-786-0011 (料金は一泊朝食付きバスは共同で￥5,500 程です。 http://ss.nrifs.affrc.go.jp/intro/hotel.html を参照してください)。